### v1

# of our esteemed colleague Professor

Abstract.
A smooth map between manifolds is said to be image simple if its restriction to its singular point set is a topological embedding. We study the parity of the number of connected components of the singular point set for image simple fold maps from a closed manifold of dimension $\geq 2$ to a surface. It is known that this parity is a homotopy invariant when the source manifold is of even dimension and the target surface is orientable. In this paper we show that for an arbitrary image simple fold map from a closed odd-dimensional manifold of dimension $\geq 3$ to a (possibly non-orientable) surface, this parity is not a homotopy invariant. We give two constructive proofs: one uses techniques from open book decompositions and round fold maps, and the other uses allowable moves. The constructed homotopies add one new component to the singular point set. We also give a new proof of the homotopy invariance of the parity of the number of connected components of the singular point set for image simple fold maps from a closed even-dimensional manifold to an orientable surface together with an example showing that this does not hold for maps into non-orientable surfaces in general.
Key words and phrases:
fold map, open book decomposition, round fold map, image simple map, parity of the number of components of singular point set
2020 Mathematics Subject Classification:
58K30; 58K65, 57R45
1. Introduction
Given a smooth map $f\colon M\to N$ between manifolds of dimensions $m$ and $n$, respectively, with $m\geq n$, we say that $x\in M$ is a singular point if the differential $df_{x}\colon T_{x}M\to T_{f(x)}N$ does not have the maximal rank $n$. It is known that when $f$ is a generic smooth map into a surface $N=S$, the set $\Sigma(f)$ of its singular points consists of folds and cusps (defined in Section 2), see [16, 17]. More specifically, for a given closed manifold $M$ of $\dim{M}\geq 2$ and a surface $S$, the set of smooth maps $f\colon M\to S$ which have only folds and cusps as their singular points is open and dense in the space $C^{\infty}(M,S)$ of smooth maps of $M$ into $S$ endowed with the Whitney $C^{\infty}$ topology (for example, see [5]). In this sense, generic maps to surfaces can be considered as generalizations of Morse functions on manifolds which are generic maps into the $1$-dimensional manifold $\mathbb{R}$. Therefore, it is expected that the topology of the source manifold $M$ is somehow related to various properties of generic maps $M\to S$ on that manifold.
In this paper, we focus on generic maps to surfaces without cusp singularities. A smooth map $f\colon M\to S$ that has only folds as its singularities is called a fold map. Note that if $M$ is a closed manifold, then the singular point set $\Sigma(f)$ is a finite disjoint union of circles in $M$, and the restriction $f|_{\Sigma(f)}$ is an immersion. Among fold maps, we are mainly interested in those whose singular point sets have well-behaved images. We say that a smooth map $f$ between manifolds is image simple if the restriction of $f$ to the set $\Sigma(f)$ of its singular points is a topological embedding. In other words, the image $f(\Sigma(f))$ of the singular point set has no self-intersections in the target manifold. Image simple smooth generic maps have recently attracted some detailed studies (see, for example, [1, 2, 8, 9, 10, 13, 15]).
Thus, our main objective is to study certain topological properties of image simple fold maps from a closed manifold $M$ of dimension $m\geq 2$ to a surface $S$. The image $f(\Sigma(f))$ of the singular point set $\Sigma(f)$ of such a map is a finite disjoint union of embedded closed curves in $S$. The main aim of the paper is to study one of the basic characteristics of an image simple fold map $f$, namely the parity of the number $|\Sigma(f)|$ of components of the singular point set, and specifically, whether this parity is a homotopy invariant of $f$ or not.
This question is a generalization of the first of the two Takase problems formulated in [13, Problem 3.2]. Takase considered image simple generic maps $f\colon S^{3}\to S^{2}$ without definite fold singular points. (For the definition of a definite fold singularity, refer to Section 2.) The first Takase problem asks if the number $|\Sigma(f)|$ of components of $\Sigma(f)$ is congruent modulo two to $d+1$, where $d\in\mathbb{Z}\cong\pi_{3}(S^{2})$ is the homotopy class represented by $f$. The second problem asks if the minimum number of components of $\Sigma(f)$ is equal to $|d|+1$.
The second author [14] later answered the first Takase problem in the negative by constructing two homotopic image simple fold maps $f$ and $g$ without definite folds such that $|\Sigma(f)|\not\equiv\#|\Sigma(g)|\pmod{2}$, see Section 4 for details. Thus, in the case of image simple fold maps of manifolds of dimension $3$ to surfaces, the parity of the number of fold components is not a homotopy invariant. It is worth mentioning that the second author constructed an explicit homotopy $F\colon M\times[0,1]\to S\times[0,1]$ between $f$ and $g$. The singular point set of a generic homotopy between fold maps to a surface is, in general, a smooth surface properly embedded in $M\times[0,1]$, and we see that for the specific homotopy $F$ constructed in [14], the surface $\Sigma(F)$ of the singular points of $F$ is the disjoint union of $\Sigma(f)\times[0,1]$ and a Möbius band. Furthermore, the restriction of $F$ to the Möbius band has a triple self-intersection point.
The mentioned results raised the broader question of identifying the precise conditions under which the parity of the number of fold components is a homotopy invariant. It was partially answered by the first author and Kahmeyer in [7], who proved that for an image simple fold map $f\colon M\to S$ of a closed manifold $M$ of dimension $\geq 2$ to a surface $S$, the parity of the number of connected components of $\Sigma(f)$ does not change under homotopy $F\colon M\times[0,1]\to S\times[0,1]$ if any of the following conditions holds:
  1. (1)
    the dimension of $M$ is even and $S$ is orientable (refer to Theorem 5.1 of the present paper as well),
  2. (2)
    the singular point set $\Sigma(F)$ of the homotopy is an orientable surface, or
  3. (3)
    the restriction $F|_{\Sigma(F)}$ of homotopy to its singular point set does not have triple self-intersection points, and either $M$ is a closed orientable manifold of dimension $3$ and $S$ is orientable, or $M$ is a closed manifold of odd dimension $m>2$ and $S$ is $\mathbb{R}^{2}$ or $S^{2}$.
In the present paper we will study the cases excluded in the above result (1): when the dimension of $M$ is odd and is at least $3$ and $S$ is an arbitrary surface, and when the dimension of $M$ is even and $S$ is a non-orientable surface. Our first main result shows that in the odd-dimensional case, the parity is not a homotopy invariant as follows.
Theorem 1.1.
Let $M$ be a closed odd-dimensional manifold with $m=\dim M\geq 3$, and $S$ be a possibly non-orientable surface. Then, for every image simple fold map $f\colon M\to S$, there exists a homotopy $F$ connecting $f$ to another image simple fold map $g$ such that $|\Sigma(f)|$ and $|\Sigma(g)|$ have different parities. Furthermore, the homotopy $F$ can be chosen so that it is supported in an arbitrarily small $m$-disc in $M$.
In fact, we will construct an explicit homotopy $F\colon M\times[0,1]\to S\times[0,1]$ satisfying the conclusion of Theorem 1.1. Its singular point set $\Sigma(F)$ contains a Möbius band component, and $F|_{\Sigma(F)}$ has a triple self-intersection point. This shows that conditions (2) and (3) above are necessary in general. On other other hand, since Theorem 1.1 permits the target surface $S$ to be non-orientable, it is not guaranteed that every homotopy satisfying the conclusion of the theorem must have a triple self-intersection point. However, if we remove the condition that the homotopy should be supported in a small disc, then, as we will see below, in the case where $M$ is even-dimensional, and both $M$ and $S$ are non-orientable, the absence of a triple self-intersection is not a necessary condition.
Our second main theorem is the following.
Theorem 1.2.
For every $n\geq 1$, there exist a closed $2n$-dimensional non-orientable manifold $M$ and image simple fold maps $f$ and $g\colon M\to\mathcal{M}$ to the open Möbius band $\mathcal{M}$ such that $f$ and $g$ are homotopic, but $|\Sigma(f)|\not\equiv\#|\Sigma(g)|\pmod{2}$.
We also construct an explicit homotopy $F$ between $f$ and $g$ in Theorem 1.2 such that the singular point set $\Sigma(F)$ is non-orientable, but that it has no triple self-intersections. This again shows that condition (2) above is necessary in general: however, condition (3) is not necessary in general.
To summarize, let $M$ be a closed manifold of dimension at least $2$, and $S$ be a surface. Let $F$ be a homotopy between two image simple fold maps $f$ and $g\colon M\to S$. Suppose that the dimension of $M$ is even.
  • If $S$ is orientable, then the parity is a homotopy invariant by [7] or by Theorem 5.1.
  • If $S$ is non-orientable, then the parity is not a homotopy invariant by Theorem 1.2.
  • If $|\Sigma(f)|\not\equiv\#|\Sigma(g)|\pmod{2}$, then $\Sigma(F)$ is non-orientable, but $F|_{\Sigma(F)}$ may or may not have triple self-intersection points by [7] and by Theorem 1.2.
Suppose that the dimension $m$ of $M$ is odd.
  • The parity of the number of fold components is not a homotopy invariant in general by [14] in the case $m=3$ and by Theorem 1.1 in the case $m\geq 3$.
  • If $\Sigma(F)$ is an orientable surface, then the parity is an invariant by [7].
Furthermore, suppose that either $M$ is a closed orientable manifold of dimension $3$ and $S$ is orientable, or $M$ is a closed manifold of odd dimension $m>2$ and $S$ is $\mathbb{R}^{2}$ or $S^{2}$.
  • If $|\Sigma(f)|\not\equiv\#|\Sigma(g)|\pmod{2}$, then $F|_{\Sigma(F)}$ has a triple self-intersection point by [7].
We do not know if the last statement remains true without the assumptions on $S$.
The paper is organized as follows. In Section 2 we prepare some basic notions necessary in the paper. In Section 3 we use open book decompositions as well as round fold maps of odd-dimensional spheres to prove Theorem 1.1 producing a homotopic modification $g$ of $f$ with $|\Sigma(g)|\equiv\#|\Sigma(f)|+1\pmod{2}$. In Section 4 we give another proof of Theorem 1.1 along the argument in [14], using allowable moves (or always-realizable moves) as listed in [15]. In Section 5, we first give a new proof of the homotopy invariance of the parity of the number of connected components of the singular point set for image simple fold maps from a closed even-dimensional manifold to an orientable surface (see Theorem 5.1). Then, we prove Theorem 1.2 by constructing an explicit example.
Throughout the paper, manifolds and maps between them are smooth of class $C^{\infty}$ unless otherwise specified. The symbol $\cong$” denotes an appropriate isomorphism between algebraic objects or a diffeomorphism between manifolds.
2. Preliminaries
Let $M$ be a smooth closed manifold of dimension $m\geq 2$, and $S$ a smooth surface.
Definition 2.1.
A singular point $p$ of a smooth map $f\colon M\to S$ that can be described by the normal form
$(t,x_{1},x_{2},\ldots,x_{m-1})\mapsto(X,Y)=(t,-x_{1}^{2}-\cdots-x_{i}^{2}+x_{i+1}^{2}+\cdots+x_{m-1}^{2})$
with respect to appropriate coordinates around $p$ and $f(p)$ is called a fold singularity (or a fold), where the integer $i$ is called the index with respect to the (Y)(-Y)-direction (see Fig. 1). When we choose the $Y$-coordinate so that $0\leq i\leq\left\lfloor(m-1)/2\right\rfloor$ holds, $i$ is called the absolute index. Note that the absolute index of a fold is well-defined and does not depend on the choice of local coordinates.
When $i=0$ or m1m-1, the singular point is called a definite fold. Otherwise, it is called an indefinite fold.
A smooth map $f\colon M\to S$ that has only folds as its singularities is called a fold map.
[…]
Figure 1. Fold and cusp singularities for $m=2$
For a smooth map $f\colon M\to S$, $\Sigma(f)$ denotes the set of singular points of $f$. If $f$ is a fold map and the source manifold $M$ is closed (compact and without boundary), then $\Sigma(f)$ is a disjoint union of finitely many circles, and $f|_{\Sigma(f)}$ is an immersion.
Definition 2.2.
A singular point of a smooth map $M\to S$ that has the normal form
$(t_{1},t_{2},x_{1},x_{2},\ldots,x_{m-2})$
$\mapsto(X,Y)=(t_{1},t_{2}^{3}+t_{1}t_{2}-x_{1}^{2}-\cdots-x_{i}^{2}+x_{i+1}^{2}+\cdots+x_{m-2}^{2})$
for some $0\leq i\leq(m-2)/2$ is called a cusp singularity (or a cusp). See Fig. 1.
Definition 2.3.
A fold map $f\colon M\to S$ of a closed manifold into a surface is said to be image simple if $f|_{\Sigma(f)}\colon\Sigma(f)\to S$ is a topological embedding. Since, for a fold map, the restriction to its singular point set is an immersion, this is equivalent to requiring that $f|_{\Sigma(f)}$ is a smooth embedding.
[…]
Let $F$ be a compact manifold of dimension $k$, and $\mu\colon F\to F$ a diffeomorphism which is the identity on the boundary $\partial F$. Then, we have a fiber bundle
$S^{1}\tilde{\times}F=[0,2\pi]\times F/\!\!\sim\,\,\to S^{1},$
where we define $(2\pi,x)\sim(0,\mu(x))$ for each $x\in F$, and where the projection map is induced by the projection to the first factor $0,2\pi]\times F\to[0,2\pi]$. Note that the boundary of this bundle is a trivial fiber bundle $S^{1}\times\partial F\to S^{1}$. Then, the space $M=S^{1}\tilde{\times}F/\!\sim$ obtained from $S^{1}\tilde{\times}F$ by identifying $S^{1}\times\{y\}$ to a point for each $y\in\partial F$ has a natural structure of a smooth closed $(k+1)$-dimensional manifold. Such a description of $M$ is called an open book decomposition with monodromy $\mu$. Furthermore, the submanifold $\{t\}\tilde{\times}F$ for each $t\in S^{1}$ is called a page and the submanifold (diffeomorphic to $\partial F$) obtained from $S^{1}\times\partial F$ by the identification is called the binding.
Suppose that there exists a smooth family $f_{t}\colon F\to[0,\infty)$ of Morse functions, $t\in S^{1}$, such that
  • $f_{t}(\partial F)=0$,
  • the critical values of $f_{t}$ are in $(0,\infty)$,
  • the critical values are all distinct, and
  • $\tilde{f}|_{\Sigma(\tilde{f})}$ is locally constant, where $\tilde{f}\colon S^{1}\times F\to[0,\infty)$ is defined by $\tilde{f}(t,x)=f_{t}(x)$.
Furthermore, we assume that $f_{2\pi}=f_{0}\circ\mu$ and hence that the family $\{f_{t}\}$ induces a smooth map $\beta\colon S^{1}\tilde{\times}F\to S^{1}\times[0,\infty)$ defined by $\beta(t,x)=(t,f_{t}(x))$. Then, the composition of $\beta$ with the natural map $S^{1}\times[0,\infty)\to\mathbb{R}^{2}$ corresponding to the polar coordinates of $\mathbb{R}^{2}$ defines a smooth map $M\to\mathbb{R}^{2}$. We see easily that the map thus constructed is a round fold map. Furthermore, the indices of its fold points with respect to the inward direction coincide with those of the corresponding critical points of the Morse functions $f_{t}$.
3. A peculiar open book decomposition of the sphere $S^{2n+1}$
For $n\geq 1$, let $P$ denote the total space of the disc bundle over $S^{n}$ associated with the tangent bundle $TS^{n}$ over $S^{n}$. Specifically, we choose $S^{n}$ to be the standard unit Riemannian sphere in $\mathbb{R}^{n+1}$, and let $P$ be the submanifold of $TS^{n}$ consisting of points $(x,v)$ with $x\in S^{n}$ and $v\in T_{x}S^{n}$ such that the length $v||$ of the vector $v$ at $x$ is at most $\pi$. The projection of this disc bundle is denoted by $\pi_{P}\colon P\to S^{n}$.
We now define a generalized Dehn twist of $P$ relative to the boundary as follows. Suppose that $\varphi$ is a self-diffeomorphism of $P$ that is the identity on the boundary. It takes a fiber $D_{x}=\pi_{P}^{-1}(x)$ over a point $x\in S^{n}$ to a properly embedded disc $\varphi(D_{x})$ in $P$. Since $\varphi$ preserves $\partial D_{x}$ point-wise, the projection $\pi_{P}|_{\varphi(D_{x})}$ takes the boundary of the disc $\varphi(D_{x})$ to the point $x$, and therefore the restriction $\pi_{P}\circ\varphi|_{D_{x}}$ defines an element in $\pi_{n}(S^{n})\cong\mathbb{Z}$. We say that the self-diffeomorphism $\varphi$ of $P$ is a generalized Dehn twist if it is the identity on the boundary, and the map $\pi_{P}\circ\varphi|_{D_{x}}$ represents a generator of the group $\pi_{n}(S^{n})\cong\mathbb{Z}$ for every fiber $D_{x}$ of $\pi_{P}$. (If we orient $S^{n}$ and the fibers of $P$ consistently, then we say that $\varphi$ is a generalized positive Dehn twist if $\pi_{P}\circ\varphi|_{D_{x}}$ represents $1\in\pi_{n}(S^{n})\cong\mathbb{Z}$.)
To construct such a diffeomorphism $\varphi$ explicitly, let $\gamma\colon P\to P$ denote the smooth map that takes a tangent vector $v$ at a point $x\in S^{n}$ to the point $(\ell(1),\ell^{\prime}(1))\in P$, where $\ell$ is the unique geodesic $\mathbb{R}\to S^{n}$ such that $\ell(0)=x$, $\ell^{\prime}(0)=v$, and $\ell^{\prime}(t)\in T_{\ell(t)}S^{n}$ denotes the velocity vector of $\ell$ at $t\in\mathbb{R}$. We see easily that $\gamma$ is a diffeomorphism, since one can construct its inverse by sending $(x,v)$ to $(\ell(-1),\ell^{\prime}(-1))$.
Lemma 3.1.
The diffeomorphism $\gamma\colon P\to P$ is isotopic to the identity through diffeomorphisms which may not fix the boundary point-wise.
Proof.
For $v\in T_{x}S^{n}$ with $v||\leq\pi$ and $t\in[0,1]$, let $\gamma_{t}(x,v)$ be the point $(\ell(t),\ell^{\prime}(t))\in P$, where $\ell$ is the unique geodesic as described above. We see easily that this defines a well-defined smooth map $\gamma_{t}\colon P\to P$. By an argument similar to the above, we see that this defines a smooth $1$-parameter family of diffeomorphisms. Furthermore, we see that $\gamma_{0}$ is the identify of $P$ and that $\gamma_{1}=\gamma$. This completes the proof. ∎
Now, we define the diffeomorphism $\varphi$ by $\varphi=\tau\circ\gamma$, where $\tau$ is the involution that takes a tangent vector $v$ at $x\in S^{n}\subset\mathbb{R}^{n+1}$ to the tangent vector v-v at x-x. Here, we are identifying the tangent space $T_{x}S^{n}$ with $\{v\in\mathbb{R}^{n+1}\,:\,x\cdot v=0\}$, where $\cdot$” denotes the usual Euclidean inner product. It immediately follows that the diffeomorphism $\varphi$ is the identity on the boundary of $P$, and that $\pi_{P}\circ\varphi|_{D_{x}}$ represents a generator of the group $\pi_{n}(S^{n})\cong\mathbb{Z}$ for every fiber $D_{x}$ of $\pi_{P}$.
We also construct a Morse function $f\colon P\to[-1,2]$ with exactly two critical points $p$ and $q$ of indices 0 and $n$, respectively, such that f(p)=1f(p)=-1, $f(q)=1$ and $f(\partial P)=2$. Namely, let $\xi\colon S^{n}\to[-1,1]$ be a standard Morse function on $S^{n}$ with two critical points $p$ and $q$ of indices 0 and $n$, respectively, and $\eta$ a Morse–Bott function on $P$ that assigns to each vector $v$ the square $v||^{2}$ of its length. Then $f=\xi\circ\pi_{P}+(2-\xi\circ\pi_{P})\cdot\eta/\pi^{2}$ is a desired Morse function on $P$. Note that the critical points $p$ and $q\in P$ of $f$ are points in the zero section $S^{n}$ of $P$. The closure of the descending disc of the critical point $q$ coincides with the zero section $S^{n}$ of $P$, while the ascending disc of $q$ is the fiber $D_{q}=\pi_{P}^{-1}(q)$, provided that we choose the gradient-like vector field of $f$ appropriately.
Let $DP=P_{R}\cup P_{L}$ denote the (untwisted) double of the manifold $P$, formed from the disjoint union of $P_{R}$ and $P_{L}$ by identifying their boundaries by the identity map. Then, $\varphi^{-1}(D_{q})$ in $P_{R}$ together with $D_{q}$ in $P_{L}$ form a sphere $S_{1}$. A copy of $D_{q}$ in $P_{R}$ and a copy of $D_{q}$ in $P_{L}$ also form a sphere $S_{2}$.
Lemma 3.2.
The sphere $S_{1}$ is isotopic, in $DP$, to a sphere that intersects $S_{2}$ transversely at a unique point.
Proof.
The sphere $S_{1}$ is isotopic to the sphere $S_{1}^{\prime}$ which is the union of a copy of $\varphi^{-1}(D_{p})$ in $P_{R}$ and a copy of $D_{p}$ in $P_{L}$, since $D_{p}$ and $D_{q}$ are isotopic through fibers. Then, we see that $S_{1}^{\prime}\cap S_{2}$ consists of $\varphi^{-1}(D_{p})\cap D_{q}$ in $P_{R}$, since $D_{p}$ and $D_{q}$ are disjoint in $P_{L}$. Since $\varphi=\tau\circ\gamma$ and $\tau^{-1}(D_{p})=D_{q}$, we have only to show that $D_{q}$ and $\gamma(D_{q})$ intersect transversely at $(q,0)$, the point of the zero section on the fiber $D_{q}$. This is obvious by the very definition of $\gamma$. This completes the proof. ∎
Let $X$ denote the closed manifold of dimension $2n+1$ with an open book decomposition given by the page $P$ and monodromy $\varphi$. We will give the description of $X$ as in [6]. The manifold $X$ can be constructed from $0,2\pi]\times P$ by performing a “vertical” identification $V$ and a “horizontal” identification $H$, where
  • $V$ identifies $(2\pi,x)$ with $(0,\varphi(x))$ for every $x\in P$, and
  • $H$ identifies $(t,y)$ with $(s,y)$ for every $y\in\partial P$ and all $t,s\in[0,2\pi]$.
Let $X_{+}$ and $X_{-}$ denote the subspaces of $X$ consisting of the equivalence classes of points in $0,\pi]\times P$ and $\pi,2\pi]\times P$, respectively. Then $X$ is the union of $X_{+}$ and $X_{-}$. The interiors of $X_{+}$ and $X_{-}$ are disjoint, while the boundaries $\partial X_{+}$ and $\partial X_{-}$ are diffeomorphic to the double $DP$ of the manifold $P$. We write $\partial X_{\pm}=P^{\pm}_{R}\cup P^{\pm}_{L}$, where $P^{+}_{R}\subset X_{+}$ stands for $\{0\}\times P$, $P^{+}_{L}$ and $P^{-}_{L}$ stand for $\{\pi\}\times P$, and $P^{-}_{R}$ stands for $\{2\pi\}\times P$. Thus, the manifold $X$ is obtained from the disjoint union of $X_{+}$ and $X_{-}$ by identifying $P^{+}_{L}$ with $P^{-}_{L}$ by means of the identity map, and by identifying $P^{-}_{R}$ with $P^{+}_{R}$ by means of $\varphi\colon P_{R}^{-}\to P_{R}^{+}$ (see Fig. 2).
[…]
Figure 2. Open book decomposition of $X$
The page $P$ has a handle decomposition that consists of a handle of index 0, and a handle $h^{n}$ of index $n$. Therefore, the manifold $X$ admits a handle decomposition with a handle of index 0, a handle of index $n$, a handle of index $n+1$ and a handle of index $2n+1$. Indeed, the manifold $X_{-}$ admits a handle decomposition with a handle of index 0 and a handle of index $n$, while $X_{+}$ admits a handle decomposition with a handle of index $n+1$ and a handle of index $2n+1$. Thus, the manifold $X$ admits a handle decomposition as described.
The attaching sphere of the $(n+1)$-handle in $X$ is the union of the cocore disc $coc(h^{n})_{L}$ of the $n$-handle $h^{n}$ in $P_{L}^{-}\cong P$ and the copy of its image $\varphi^{-1}(coc(h^{n})_{R})$ in $P_{R}^{-}\cong P$. By Lemma 3.2, the attaching sphere $coc(h^{n})_{L}\cup\varphi^{-1}(coc(h^{n})_{R})$ of the $(n+1)$-handle in $X$ is isotopic to a sphere that intersects the belt sphere $coc(h^{n})_{R}\cup coc(h^{n})_{L}$ of the $n$-handle in $X$ at a unique point transversely, and therefore the $n$-handle and the $(n+1)$-handle in $X$ are canceling. We conclude that the manifold $X$ is obtained by gluing two smooth $(2n+1)$-dimensional discs. Therefore, $X$ is a smooth manifold $\Sigma^{2n+1}$ homotopy equivalent to a sphere. In particular, for $n=1,2$, $X$ is diffeomorphic to the standard sphere $S^{2n+1}$.
We claim that the monodromy $\varphi$ can be modified, if necessary, in a neighborhood $U$ of a point, say $x_{0}\in\mathrm{Int}\,P=\mathrm{Int}\,P_{R}^{-}$, so that the resulting manifold $X$ is diffeomorphic to the standard $(2n+1)$-sphere, provided that $n\geq 3$. Furthermore, we may choose $U$ such that its closure is contained in $P_{R}^{-}\setminus Q$, where $Q$ is the union of the zero section $S^{n}$ of $P_{R}^{-}$ and $\varphi^{-1}(D_{q})$. Indeed, assume that $X$ is an exotic sphere $\Sigma^{2n+1}$. It is obtained by gluing two discs $D^{2n+1}_{-}$ and $D^{2n+1}_{+}$ by means of a diffeomorphism $\phi\colon S^{2n}\to S^{2n}$. By modifying $\phi$ by isotopy if necessary, we may assume that $\phi$ is the identity map except on a small disc neighborhood $B^{2n}$ of a point. Let $\phi_{B}\colon B^{2n}\to B^{2n}$ be the restriction of $\phi$ to $B^{2n}$. Let us embed $B^{2n}$ to a small neighborhood $U$ of a point $x_{0}$ in the interior of $P_{R}^{-}$ such that the closure of $U$ is contained in $P_{R}^{-}\setminus Q$. Then, by modifying $\varphi\colon P_{R}^{-}\to P_{R}^{+}$ on $B^{2n}$ using $\phi_{B}^{-1}$, we get a manifold diffeomorphic to $X\sharp(-\Sigma^{2n+1})\cong S^{2n+1}$, where $\Sigma^{2n+1}$ is the manifold $\Sigma^{2n+1}$ with the orientation reversed (see [3]).
Recall that $P$ admits a Morse function $f$ with a unique critical point $p$ of index 0 and a unique critical point $q$ of index $n$. Now, we construct a smooth family of functions on $P$:
$f_{t}\colon P\longrightarrow\mathbb{R}$
parametrized by $t\in[0,2\pi]$ such that $f_{0}=f$, $f_{2\pi}=f\circ\varphi$, $f_{t}$ is a Morse function with a unique critical point of index 0 with value 1-1, a unique critical point of index $n$ with value $1$, and $f_{t}(\partial P)=2$ for all $t\in[0,2\pi]$. To this end we will construct an ambient isotopy $g_{t}$ of $P$ parametrized by $t\in[0,2\pi]$ such that $f_{t}=f\circ g_{t}$ is a desired family of Morse functions.
To construct an ambient isotopy $g_{t}$, we first consider an isotopy parametrized by $t\in[0,\pi]$ of the ascending disc $\varphi^{-1}(D_{q})=\gamma^{-1}(\tau^{-1}(D_{q}))$ of $f\circ\varphi$ to the ascending disc $D_{q}$ of $f$. Such an isotopy can be constructed by using the ambient isotopy $\gamma_{t}$, $t\in[0,1]$, followed by an isotopy taking $\tau^{-1}(D_{q})$ to $D_{q}$, where $\gamma_{t}$, $t\in[0,1]$, is the ambient isotopy of $P$ constructed in the proof of Lemma 3.1. Furthermore, we may choose an ambient isotopy $\tau_{t}$ of $P$ that takes $\tau^{-1}(D_{q})$ to $D_{q}$ so that it consists of diffeomorphisms of $P$ induced by rotations of $S^{n}$ about any axis perpendicular to the line $\overline{pq}$ passing through $p$ and $q$ in $\mathbb{R}^{n+1}$. Here, note that when $n$ is even, the isotopy thus constructed reverses the orientations of $\varphi^{-1}(D_{q})$ and $D_{q}$.
Finally, we can modify the isotopies $\gamma_{t}$ and $\tau_{t}$ slightly so that their concatenation defines a smooth ambient isotopy $g_{t}$ of $P$ parametrized by $t\in[0,\pi]$.
Note that the Morse function $f\circ\varphi\colon P\to[-1,2]$ has critical points $q$ and $p$ of indices 0 and $n$, respectively. Then, we have $f\circ g_{\pi}=f\circ\varphi$ over the ascending disc $\varphi^{-1}(D_{q})$ of $p$ as well as the closure of the descending disc $S^{n}$ of $p$. Indeed, since $\gamma_{t}$ and $\tau_{t}$ preserve the lengths of the vectors in $P$, we conclude that $f\circ g_{\pi}=f\circ\varphi$ over $\varphi^{-1}(D_{q})$. On the other hand, $\gamma$ is the identity on $S^{n}$, while $\tau$ takes every level set of $f|_{S^{n}}$ to a level set. Therefore, we have $f\circ g_{\pi}=f\circ\varphi$ on $S^{n}$. Note that we have $g_{\pi}(Q)=\varphi(Q)$ even though $g_{\pi}$ may not be isotopic to $\varphi$ on $Q$.
We observe that $P\setminus Q$ is diffeomorphic to $(\partial P\setminus\varphi^{-1}(\partial D_{q}))\times(-1,2]$ under the diffeomorphism that takes a point $x\in P\setminus Q$ to the pair $(y,f\circ g_{\pi}(x))$, where $y$ is the unique point on $\partial P\setminus\varphi^{-1}(\partial D_{q})$ that lies on the gradient curve of $f\circ g_{\pi}$ passing through $x$. There is an isotopy $g_{t}$ of $P\setminus Q$ parametrized by $t\in[\pi,2\pi]$ that takes a point $x$ with coordinates $(y,f\circ g_{\pi}(x))$ to the point with coordinates $(y,f\circ\varphi(x))$. Such an isotopy extends over $Q$, since $f\circ g_{\pi}=f\circ\varphi$ over $Q$, and we can also modify the two isotopies $g_{t}$ for $t\in[0,\pi]$ and $t\in[\pi,2\pi]$ to produce a smooth isotopy of $P$ parametrized by $t\in[0,2\pi]$.
By replacing $f_{t}$ by $f_{t}-2$, we arrange the Morse functions $f_{t}$ so that $f_{t}(\partial P)=0$, $t\in[0,2\pi]$. Define the map
$0,2\pi]\times P\longrightarrow\mathbb{R}^{2}$
by $\rho(t,x)=-f_{t}(x)$ and $\theta(t,x)=t$, where $(\theta,\rho)$ are the polar coordinates on $\mathbb{R}^{2}$. The map thus constructed defines a well-defined smooth map $f_{X}\colon X\to\mathbb{R}^{2}$. This is an image simple round fold map with a unique fold circle of index $2n$, and a unique fold circle of index $n$ with respect to the inward direction.
Now, let $f\colon M\to S$ be an arbitrary image simple fold map into a (possibly non-orientable) surface $S$. We first consider the disjoint union $f\cup f_{X}\colon M\cup X\to S$, where $f_{X}(X)$ is contained in a small $2$-disc in $f(M)\setminus f(\Sigma(f))$. We apply the birth move (see [15, Fig. 4]) to create a wrinkle consisting of two cusps together with fold arcs of absolute indices 0 and $1$ in such a way that the image sits just next to $f_{X}(\Sigma(f_{X}))$. (For this and the following, see Fig. 3. The integers attached to some curves indicate the corresponding indices of the relevant folds with respect to the normal directions indicated by arrows.) Then, we construct an image simple fold map $M\#X\cong M\to S$ by the connected sum operation along definite folds, as described in [12, Lemma 5.4], in such a way that the image of the singular point set is as depicted in the upper right figure of Fig. 3. Note that the fold index $n$ does not change even if we change the normal direction, since the dimension of the source manifold is equal to $2n+1$. We apply the move of type II twice so that the image of the fold circle of index $n$ gets out of the wrinkle. This is possible as long as n<2n1n<2n-1, i.e. $n\geq 2$ (see [15, Fig. 5]). Finally, we can eliminate the wrinkle, which is possible, since one of the fold arcs has absolute index 0 (see [4]). The resulting fold map $g\colon M\to S$ is image simple and we have $|\Sigma(g)|=\#|\Sigma(f)|+1$.
This completes the proof of Theorem 1.1 for $n\geq 2$.
[…]
Figure 3. Sequence of the singular point set images
Remark 3.3.
(1) We do not know if a similar argument works for $n=1$. We will give another argument to cover this case in Section 4. On the other hand, if we merge the two cusp points in the upper right figure of Fig. 3 using a path whose image is embedded in the left hand side of the wrinkle, we get an image simple fold map $h\colon M\to S$ homotopic to $f$ such that $|\Sigma(h)|=\#|\Sigma(f)|+3$. This proves Theorem 1.1 for the case of $n=1$.
(2) We see easily that $f$ and $g$ are homotopic as continuous maps. However, we do not know how to construct explicit homotopy between $f$ and $g$ thus constructed.
4. Allowable moves approach
Let us give another proof to Theorem 1.1 by an approach similar to that in [14]. For the moves that we use in this section, the reader is referred to [15].
Suppose $n\geq 2$. Let $f\colon M\to S$ be an arbitrary image simple fold map of a closed $(2n+1)$-dimensional manifold into a (possibly non-orientable) surface $S$. We consider a small $2$-disc $\Delta$ contained in $f(M)\setminus f(\Sigma(f))$, and we modify the map $f$ on a connected component of $f^{-1}(\Delta)$, following the procedure as described in Fig. 4. These figures depict the singular value sets of generic maps inside $\Delta$. The integers attached to some curves indicate the corresponding indices of the relevant folds with respect to the normal directions indicated by arrows. The lower left figure depicts the singular value set inside the $2$-disc $\Delta$, i.e. the empty set. Then, we create three wrinkles, using the birth move three times (see [15, Fig. 4]). Then, by three flip moves, we turn the wrinkles into three bow ties with involved folds of indices n1n-1, $n$ and $n+1$ (see [15, Fig. 4]). Note that the index n1n-1 turns into $n+1$ if we reverse the normal direction. Then, by three cusp merges we merge the three bow ties as in the upper right figure. This is possible, since we are modifying the maps only on a connected component of $f^{-1}(\Delta)$ so that the relevant fibers over the points on the paths between the images of the cusps that are merged are connected. As the three fold curves all have indices $n\geq 2$, we can perform the move of type III as in the lower right figure (see [15, Fig. 6]). Finally, we need to perform three unflips. This is possible, since the involved folds are of indices $n$ and $n+1$ with $2\leq n\leq 2n-1$ (see [15, Fig. 4]).
[…]
Figure 4. Sequence of moves
The above argument may not work directly for $n=1$. However, as noted in [14], by merging the three pairs of cusps with appropriate “twists”, we can guarantee that the move of type III can be applied.
This completes the proof of Theorem 1.1.
This approach allows us to determine the singular point set of the constructed homotopy as follows. As we perform three births, we have three $2$-discs as the singular point set in the first three steps. In the fourth step, we merge the $2$-discs by three bands in such a way that the resulting compact surface has exactly one boundary circle whose image is depicted in the upper right figure of Fig. 4. Therefore, the singular point set for the first four steps must be a Möbius band. The rest of the homotopy does not change the topological type of the singular point set component. Therefore, the resulting singular point set of the homotopy is a Möbius band, except for the cylinder part corresponding to the fold circles of the original image simple fold map $f$.
So, the singular point set of the entire homotopy $F\colon M\times[0,1]\to S\times[0,1]$ is diffeomorphic to the disjoint union of $|\Sigma(f)|$ copies of the cylinder together with a Möbius band. Note also that $F|_{\Sigma(F)}$ has a triple self-intersection point.
This is consistent with the results obtained in [7] when the target surface $S$ is orientable and $M$ is a closed orientable manifold of dimension $3$, or $S=\mathbb{R}^{2}$ or $S^{2}$ and $M$ is a closed orientable manifold of odd dimension $m$ with $m>2$.
Note that in Theorem 1.1, the target surface $S$ is possibly non-orientable so that we do not know if for every homotopy $F$ as in Theorem 1.1, $F|_{\Sigma(F)}$ has a triple self-intersection point.
Remark 4.1.
Let us give a short remark concerning the first Takase problem formulated in [13, Problem 3.2] mentioned in Section 1. The map $f\colon S^{3}\to\mathbb{R}^{2}$ constructed in [14] does not give a direct negative answer to the problem. However, if we use the procedure as described above, we get a counter example for every element $n$ of $\pi_{3}(S^{2})\cong\mathbb{Z}$, since the newly created fold circle is indefinite. Therefore, the number modulo two of fold components of such image simple fold maps without definite fold points is not a homotopy invariant in general.
5. Even dimensional case
The following has been obtained in [7, Theorem 1.1] by using cumulative winding number in $\frac{1}{2}\mathbb{Z}$. Here, we give another proof, which seems to be more elementary in nature.
Theorem 5.1.
Let $M$ be an even-dimensional closed manifold of dimension $m\geq 2$ and $S$ a connected orientable surface. Then, for image simple fold maps $f\colon M\to S$, the parity of the number of connected components $|\Sigma(f)|$ of the singular point set is a homotopy invariant of $f$.
Proof.
Consider an image simple fold map $f\colon M\to S$. Then, $f(\Sigma(f))$ is a finite disjoint union of embedded closed curves in $S$, and we denote by $R_{1},R_{2},\ldots,R_{s}$ the connected components of $S\setminus f(\Sigma(f))$. We also denote by $\chi_{i}\in\mathbb{Z}$ the Euler characteristic of the fiber over a point in $R_{i}$, $i=1,2,\ldots,s$. When $R_{i}$ is not contained in the image of $f$, we set $\chi_{i}=0$. Note that the regular fibers of $f$ are of even dimension m2m-2 and that the parity of their Euler characteristics is a homotopy invariant of $f$. We denote it by $\chi_{f}\in\mathbb{Z}/2\mathbb{Z}$. Set $r_{f}=0\in\mathbb{Z}/4\mathbb{Z}$ if $\chi_{f}=0$ and set $r_{f}=1\in\mathbb{Z}/4\mathbb{Z}$ otherwise.
[…]
First suppose that $S$ is a closed connected orientable surface. Then, we have
$\chi(M)$ $=$ $\sum_{i=1}^{s}\chi_{i}\chi(R_{i})$
$\equiv$ $r_{f}\chi(\tilde{R}_{0})+(r_{f}+2)\chi(\tilde{R}_{1})\pmod{4}$
$\equiv$ $r_{f}\chi(\tilde{R}_{0})+(r_{f}+2)(\chi(S)-\chi(\tilde{R}_{0})\pmod{4}$
$\equiv$ $2\chi(\tilde{R}_{0})+(r_{f}+2)\chi(S)\pmod{4}$.
As the closure of $\tilde{R}_{0}$ is a compact orientable surface with boundary $f(\Sigma(f))$, we see that $2\chi(\tilde{R}_{0})\equiv 2\#|\Sigma(f)|\pmod{4}$. Since $\chi(M)$ and $(r_{f}+2)\chi(S)\pmod{4}$ are homotopy invariants of $f$, we see that the parity of $|\Sigma(f)|$ is a homotopy invariant of $f$.
In the general case where $S$ may not be closed, for two homotopic image simple fold maps $f$ and $g\colon M\to S$, we consider an appropriate compact sub-surface $\bar{S}$ of $S$ that contains the image of the homotopy between $f$ and $g$. Let $\widehat{S}$ be the closed orientable surface obtained by attaching $2$-discs to the boundary circles of $\bar{S}$. Then, for image simple fold maps $f$ and $g\colon M\to\widehat{S}$, the above argument implies that $|\Sigma(f)|\equiv\#|\Sigma(g)|\pmod{2}$.
This completes the proof. ∎
As a corollary to the above proof, we have the following.
Corollary 5.2.
Let $M$ be an even-dimensional closed manifold of dimension $m\geq 2$ and $S$ a connected orientable surface. Suppose that $M$ admits an image simple fold map $f\colon M\to S$. Then $\chi(M)$ is even. If $S$ is not closed, then we have
$|\Sigma(f)|\equiv\frac{\chi(M)}{2}\pmod{2}$.
If $S$ is closed, then $\chi(S)$ is also even, and we have
$|\Sigma(f)|\equiv\frac{\chi(M)}{2}+\chi_{f}\frac{\chi(S)}{2}\pmod{2}$,
where $\chi_{f}=0\in\mathbb{Z}/2\mathbb{Z}$ if the regular fibers of $f$ have even Euler characteristics, and $\chi_{f}=1\in\mathbb{Z}/2\mathbb{Z}$ otherwise.
In the case where the target surface is non-orientable, Theorem 5.1 does not hold in general. Let us construct an explicit example as follows.
Example 5.3.
Let $w$ and $z\colon S^{1}\to\mathbb{R}$ be two Morse functions such that (refer to Fig. 5)
  • $w$ has exactly two critical points $a_{1}$ and $a_{2}$ of indices 0 and $1$, respectively, with $w(a_{1})=-1$ and $w(a_{2})=1$,
  • $z$ has exactly four critical points $b_{1},b_{2},b_{3}$ and $b_{4}$ of indices $0,0,1$ and $1$, respectively, with $z(b_{1})=-1$, $z(b_{2})=-1/2$, $z(b_{3})=1/2$ and $z(b_{4})=1$,
  • there exists an orientation reversing involution $\sigma\colon S^{1}\to S^{1}$ such that $w\circ\sigma=-w$, $\sigma(a_{1})=a_{2}$ and $\sigma(a_{2})=a_{1}$,
  • there exists an orientation reversing involution $\tau\colon S^{1}\to S^{1}$ such that $z\circ\tau=-z$, $\tau(b_{1})=b_{4}$, $\tau(b_{2})=b_{3}$, $\tau(b_{3})=b_{2}$ and $\tau(b_{4})=b_{1}$.
[…]
Figure 5. Two Morse functions on $S^{1}$
Then, $w$ induces a fold map
$\mathrm{id}_{[0,1]}\times w\colon[0,1]\times S^{1}\to[0,1]\times\mathbb{R},$
which further induces a fold map
$\tilde{w}=\mathrm{id}_{S^{1}}\tilde{\times}w\colon S^{1}\tilde{\times}S^{1}\to S^{1}\tilde{\times}\mathbb{R}$,
where $\mathrm{id}_{[0,1]}$ is the identity map of the interval $[0,1]$, $S^{1}\tilde{\times}S^{1}$ is obtained from $0,1]\times S^{1}$ by identifying $(1,x)$ with $(0,\sigma(x))$ for each $x\in S^{1}$, and $S^{1}\tilde{\times}\mathbb{R}$ is obtained from $0,1]\times\mathbb{R}$ by identifying $(0,t)$ with (1,t)(1,-t) for each $t\in\mathbb{R}$. Similarly, we get a fold map
$\tilde{z}=\mathrm{id}_{S^{1}}\tilde{\times}z\colon S^{1}\tilde{\times}S^{1}\to S^{1}\tilde{\times}\mathbb{R}$.
We see that $\tilde{w}$ and $\tilde{z}$ are image simple fold maps from the Klein bottle to the open Möbius band $\mathcal{M}$ and that $\sharp|\Sigma(\tilde{w})|=1$, while $\sharp|\Sigma(\tilde{z})|=2$. On the other hand, since $w$ and $z$ are null-homotopic, we see that $\tilde{w}$ and $\tilde{z}$ are homotopic. Hence, the parity of the number of singular point set components is not a homotopy invariant for the image simple fold maps $\tilde{w}$ and $\tilde{z}$.
The above construction can be generalized to produce functions $w$ and $z\colon S^{2n-1}\to\mathbb{R}$, $n\geq 1$, that lead to image simple fold maps $\tilde{w}=\mathop{\mathrm{id}}_{S^{1}}\tilde{\times}w$ and $\tilde{z}=\mathop{\mathrm{id}}_{S^{1}}\tilde{\times}z$ of the $2n$-dimensional Klein bottle $S^{1}\tilde{\times}S^{2n-1}$ into the open Möbius band $\mathcal{M}=S^{1}\tilde{\times}\mathbb{R}$ such that $\tilde{w}$ has a single fold component, while $\tilde{z}$ has two fold components, as follows.
The function $w$ is the standard height function on the unit sphere $S^{2n-1}\subset\mathbb{R}^{2n}$, where $n\geq 1$, i.e., $w$ is the restriction of the last coordinate function $x_{2n}$ in $\mathbb{R}^{2n}$ to $S^{2n-1}$. Let $\sigma\colon S^{2n-1}\to S^{2n-1}$ be the orientation reversing involution defined by the reflection with respect to the coordinate hyperplane $x_{2n}=0$. Note that $w\circ\sigma=-w$.
To define the function $z$, let us observe that $S^{2n-1}$ can be obtained from the disjoint union of two solid tori $S^{n-1}\times D^{n}$ and $D^{n}\times S^{n-1}$ by identifying their boundaries via the identity map:
$\partial(S^{n-1}\times D^{n})=S^{n-1}\times S^{n-1}\stackrel{{\scriptstyle\mathop{\mathrm{id}}}}{{\longrightarrow}}S^{n-1}\times S^{n-1}=\partial(D^{n}\times S^{n-1}).$
Here, in order to get $S^{2n-1}$ which is oriented, we need to orient $S^{n-1}\times D^{n}$ and $D^{n}\times S^{n-1}$ in such a way that the identity map as above should be orientation reversing.
Since the solid torus $S^{n-1}\times D^{n}$ is a union of a 0-handle and an (n1)(n-1)-handle, this description yields a handle decomposition of $S^{2n-1}$ that consists of four handles of indices 0, n1n-1, $n$ and $2n$ respectively. In turn, such a handle decomposition gives rise to an associated Morse function $z$ on $S^{2n-1}$. There is an involution $\tau$ on $S^{2n-1}$ that swaps the two solid tori $S^{n-1}\times D^{n}$ and $D^{n}\times S^{n-1}$. By the above remark, $\tau$ should be orientation reversing. Furthermore, we can choose $z$ so that $z\circ\tau=-z$. Then, the rest of the argument is parallel to the above argument.
This proves Theorem 1.2.
We can construct a generic homotopy between the image simple fold maps $\tilde{w}$ and $\tilde{z}\colon S^{1}\tilde{\times}S^{2n-1}\to\mathcal{M}$ as follows. We first apply the birth move (see [15, Fig. 5]) to $\tilde{w}$ in such a way that we have a wrinkle consisting of two cusps together with fold arcs of indices n1n-1 and $n$ near the center circle of $\mathcal{M}$, i.e., inside the region $R$ bounded by $)\mathcal{M}\setminus\tilde{w}(\Sigma(\tilde{w}$)). Note that if we reverse the normal direction, the indices turn to $n$ and n1n-1. Then, we merge the two cusps using a path connecting them in the source manifold whose image is embedded in $R$. This is possible for $n>1$, since the fibers over the points in $R$ are connected. For $n=1$, the fibers are not connected: however, we can use a similar cusp elimination technique as in [11]. Then, we get an image simple fold map equivalent to $\tilde{z}$. The singular point set of the homotopy thus constructed is the disjoint union of a Möbius band and a cylinder, and it has no triple self-intersections.
Acknowledgment
The authors would like to express their deep gratitude and respect, with sadness, to Professor Maria Aparecida Soares Ruas, Cidinha, for her continual encouragement for pursuing research, especially in singularity theory and topology. This paper was respectfully written responding to her kind request to write something of geometric topology.
The authors would like to thank David Gay who kindly explained to the authors how to construct generalized Dehn twists. They are also thankful to Liam Kahmeyer for discussions of the topic. The second author would also like to thank the Department of Mathematics, Kansas State University, for the hospitality during the preparation of the manuscript.
This work has been partially supported by JSPS KAKENHI Grant Numbers JP22K18267, JP23H05437.